OpenStudy (anonymous):

Find the average of the function over the given interval and all values of x in the interval for which the function equals its average value. (Round your answer to three decimal places.) f(x) = 8 cos(x), (0, π/6)

OpenStudy (anonymous):

integrate and then divide by the length of the path

OpenStudy (anonymous):

i.e. find \[\frac{6}{\pi}\int_0^{\frac{\pi}{6}} 8 \cos(x)dx\]

OpenStudy (anonymous):

\[f_{average} = \frac{1}{measure(D)}\int_D f\]

OpenStudy (anonymous):

So yeah, what satellite said.

OpenStudy (anonymous):

easy enough integral, since the antiderivative is just \[8\sin(x)\] get \[\frac{6}{\pi} \sin(\frac{\pi}{6})\] \[\frac{6}{\pi}\times \frac{1}{2}=\frac{3}{\pi}\]

OpenStudy (anonymous):

bet this is class presupposes no measure theory though.

OpenStudy (anonymous):

oh damn forgot the 8 !

OpenStudy (anonymous):

answer is \[\frac{24}{\pi}\]

OpenStudy (anonymous):

now you get to try to solve \[8\sin(x)=\frac{24}{\pi}\] have fun start with \[\sin(x)=\frac{3}{\pi}\] and then maybe graph?

OpenStudy (anonymous):

i get some arcsin stuff ugh how do i solve that

OpenStudy (anonymous):

...